OpenStudy (callisto):

Inequalities #1 Four integers a, b, c and d (where b>0) satisfy three conditions a+b=c, b+c=d and c+d=a. Find the greatest possible value of a+b+c+d

OpenStudy (shubhamsrg):

a+b =c and c+d =a so a+b+c+d = a+c => b = -d and b+c = d => c= 2d so a =3d so a+c = 5d now b>0 thus d<0 max d = just LHL of 0 so max for a+b+c+d = LHL of 0

OpenStudy (shubhamsrg):

or in layman language,,its 0.

OpenStudy (callisto):

I see!! I didn't notice that last part!!! Thanks :)

OpenStudy (shubhamsrg):

glad to help ^_^